Mark sees a crushed soda can on the ground. He kicks it fiercely with a velocity of 12m/s at projected angle of 36°. What is the highest point that the soda can reaches? Sin 36 degrees equals 0.866. 2. A 5000 kg. freight train traveling at speed of 28 m/s strikes an identical car at rest. If they stick


1
Expert's answer
2021-03-29T09:03:07-0400

Solution.

"1. v_0=12ms^{-1};"

"\alpha =36^o;"

"h_{max}-?;"

"v=0ms^{-1};"

"v_{0y}=v_0\sdot sin \alpha;" "v_{0y}=12ms^{-1}\sdot0.866=10.392ms^{-1};"

"h_{max}=\dfrac{v_y^2-v_{0y}^2}{2g};"

"h_{max}=\dfrac{0m^2s^{-2}-108m^2s^{-2}}{-2\sdot9.8ms^{-2}}=5.51m" ;

"2. m=5000kg;"

"v_1=28ms^{-1};"

"v_2=0ms^{-1};"

"m_1=m_2=m;"

"mv_1+mv_2=(m_1+m_2)v_F" ;

"v_F=\dfrac{mv_1+mv_2}{m_1+m_2};"

"v_F=\dfrac{5000kg\sdot28ms^{-1}+5000kg\sdot0ms^{-1}}{5000kg+5000kg}=14ms^{-1};"

Answer: "1. h_{max}=5.51m;\n2. v_F=14ms^{-1}."